to a man walking due east at the rate of 2km/h ,rain drops to fall vertically and when he increase his speed to 4km/h ,it appears to meet him at 25degree find real direction and speed of rain?
1
Expert's answer
2012-06-05T10:23:59-0400
Let Vm is the speed of man.
Let V is the speed and Vhor is the horizontal
component of the speed
of rain drops, A as the angle between rain's direction
and the normal.
Vhor = V * sinA.
Vhor = Vm1
V * sin25 = Vm2
V = Vm2
/ sin25 = 4 / sin25 = 9.4648 km/h
V * sinA = Vm1
A = asin(Vm1 / V) =
asin(0.2113) = 12.1991 degrees.